# AutoVP: An Automated Visual Prompting Framework and Benchmark

### v1

@@ 1 Introduction
Originating in the domain of natural language processing, prompting (Gao et al., 2021; Lester et al., 2021) has gained considerable popularity as a parameter-efficient fine-tuning approach for adapting pre-trained models to downstream tasks. Prompting’s methodology has recently been extended to the field of computer vision, where it is termed visual prompting (VP) (Bahng et al., 2022). In its simplest form, VP can be perceived as an in-domain model-reprogramming technique (Chen, 2022). More specifically, it adjusts the inputs and outputs of a pre-trained vision model to address downstream image-classification tasks, without having to make any changes to the weights or architecture of the source model’s pre-trained backbone. As such, it stands in contrast to conventional transfer-learning methods that involve complete fine-tuning, LP (i.e., involving modifications of the trainable linear layer in the penultimate layer’s output), or zero-shot learning (Radford et al., 2021). For instance, as illustrated in Figure 1, VP adds a universal trainable data frame to the target samples at the model-input stage, and then employs a mapping function – which can be either explicitly defined or implicitly learned – to associate the source and target labels at the output stage.
@@
ILM-VP ResNet18 48 IterMap 1 45.19
CLIP 30 IterMap 1 78.45
@@ 2 Background and Related Work
Traditionally, to derive task-specific machine-learning models, researchers have to train or update all model parameters. But, amid the advancement of powerful foundation models, model fine-tuning and training from scratch have both become time-consuming and parameter-inefficient approaches, usually requiring large amounts of training data and storage space. To this end, VP, also known as in-domain model reprogramming, has emerged as an effective means of obtaining machine-learning models for various domain-specific tasks (Chen, 2022). A well-developed pre-trained model from a source domain can be directly used for performing tasks in the target domain with little transformation of the target data. For example, we can use an ImageNet pre-trained model to classify medical images without modifying any of its parameters (Tsai et al., 2020). As compared to traditional approaches such as transfer learning, model fine-tuning, and training from scratch, VP is a low-complexity and model-agnostic strategy; and it is especially suitable for low-data domains.
@@
A VP framework can generally be divided into two trainable modules, one for input transformation and the other for output transformation. These are respectively placed at the input and output ends of a pre-trained model. In the case of input transformation, previous literature has proposed various ways to generate and place visual prompts. One of the most popular such approaches is to pad a frame around the target task image and then fill it with trainable additive-input perturbation (prompts) (Bahng et al., 2022; Chen et al., 2023b; Elsayed et al., 2019; Tsai et al., 2020; Wu et al., 2022). Next, since the output logits of the source pre-trained model are still in the source domain, further output transformation (e.g., LM) is required to obtain the target-domain logits. One naive way of achieving this is randomly mapping (RandMap) $m$ source labels onto the target labels. Tsai et al. (2020) found that frequency-based LM (FreqMap), which constructs its LM from the source-label prediction distribution of the target-domain data, can further improve the accuracy of downstream tasks. However, Chen et al. (2023b) argued that FreqMap lacks interpretability and that its interaction with VP is difficult to measure. To address that problem, the authors proposed iterative LM (IterMap), a transformation of FreqMap that enables it to concurrently design LM and visual prompts. Yang et al. (2023), meanwhile, proposed a semantics-based LM approach that aligns source and target classes that have similar semantic meanings. And Liao et al. (2023) utilized a prototypical verbalizer to map a mask token to downstream labels, thus providing a different perspective on LM. In this paper, we follow a similar design to Bahng et al. (2022), in which visual prompts are placed around images for input transformations, and there are four mapping methods for output transformations. Further details will be presented in Section 3.
@@
Although input prompting is commonly applied directly to the target image, researchers have also developed visual prompt tuning (Jia et al., 2022; Sohn et al., 2023), which learns prompting parameters at intermediate layers of a source model. However, this approach is primarily developed for vision transformers. In this paper, we focus on a pixel-level VP approach using a task-specific prompting model for each image-classification task. As such, our approach closely resembles real-world scenarios in which a pre-trained source model remains unmodified, and external variations are not introduced internally.
@@ 3 AutoVP Framework
In our implementation of AutoVP, we choose frame-shape prompts as the default prompting template. Hence, the visual prompt sizes $p$ represent the width of the frame, and its actual number of parameters is $2cp(h+w-2p)$, where $c$, $w$, and $h$ are color channels, width and height respectively. Although the input image size is determined by the source model, when fine-tuning to a downstream dataset from a source model, there is design freedom to resize the target images and use the remaining space for visual prompts. For instance, if the source model takes images with size $224\times 224$ as input, one can scale the target image size to $128\times 128$, resulting in the final visual prompt of size $p=(224-128)/2=48$. It was shown in Bahng et al. (2022) that the prompt size ($p$) plays a key role in VP performance. To automate the process of optimizing image resizing scale, we design a gradient-based optimization algorithm to implement the input scaling module, which is implemented using kornia.geometry.transform() from the Kornia library (Riba et al., 2020). The transform() function integrates a range of geometric transformations for 2D images into deep learning, including differentiable image scaling. In addition to image resizing, the prompt size $p$ will also scale up or down to fill the remaining space. Furthermore, to facilitate the optimization of image resizing and avoid bad local minima, we set the default image size to 128 along with three initial scales: 0.5, 1.0, and 1.5 to optimize, and the corresponding prompt sizes $p$ are 80, 48, and 16 respectively. Consequently, the input scaling module allows AutoVP to obtain the optimal image resizing scale and prompt size ($p$).
@@
[…]
@@
After applying the prompts to the resized image through the preceding stages, the prompted image is subsequently fed into the pre-trained model, which serves as a feature extractor to generate predictions in the source domain. We include four representative pre-trained models in our AutoVP framework: ResNet18 (He et al., 2016), ResNeXt101-IG (Mahajan et al., 2018), Swin-T (Liu et al., 2021), and a vision-language multi-modal model, CLIP (Radford et al., 2021). Note that in AutoVP, the weights of the pre-trained classifiers are frozen and kept unchanged. The details of the models are provided in Appendix A.1.
@@
The pre-trained models predict target data to source labels, while the last mile for VP is to map predictions on the source labels to target classes. As illustrated in Figure 1, AutoVP provides four types of output mapping and they can be generally categorized into two groups. (i) nonparametric label mapping: frequency mapping (FreqMap) and semantic mapping (SemanticMap), which are defined during the initialization of VP training and remain unchanged throughout the training process; and (ii) trainable label mapping: iterative label mapping (IterMap) and fully connected layer mapping (FullyMap). These two methods dynamically adjust the mapping based on the prompted images. In the following, we provide the overview of our four output mapping approaches, please refer to Appendix A.2 for more details.
@@
    The objective of FreqMap is to map the target label $y_{\text{t}}$ to the source label $y_{\text{s}}^{*}$, which is the most frequent label that $f_{\theta_{\text{s}}}$ classified $\tilde{\mathcal{X}_{\text{t}}}$ as. If a source class is selected as the most frequently predicted class for multiple target classes, it will be assigned to the target class that has the highest count of predictions. The general many-to-one frequency mapping algorithm is provided in Algorithm 1 in the Appendix. Moreover, random label mapping (RandMap) can be viewed as a special case of FreqMap by randomly assigning a subset of source labels to a target label.
@@ 4.1 Experimental Results
We observed that AutoVP employed FullyMap as the output transformation on most datasets. We speculate that this might have been because the linear layer has more parameters and thus allows the achievement of better results. Also, when AutoVP selected initial image scales, it had a tendency to scale up those images with relatively small prompt sizes. This allowed the VP model to allocate more attention to the image itself, leading to improved overall performance. As shown in Figure 1, when ResNet18 was used as the source model, AutoVP outperformed ILM-VP by 24.8% on average. More experimental results under this setting are provided in Appendix B.7.
@@
We also allowed AutoVP to search the optimal source model for downstream tasks. The optimal settings selected by AutoVP, and a comparison of experimental results can be found in Appendix B.7. Our experimental results show that Swin-T was the pre-trained model most frequently chosen by AutoVP as most suitable, i.e., in the cases of eight of the 12 datasets. On average, this choice resulted in 0.43% better accuracy than when CLIP was utilized as the fixed pre-trained backbone. On the DTD and Flowers102 datasets, however, Swin-T’s performance was better than CLIP’s by much more: i.e., 6.80% and 3.08%, respectively. These findings highlight how multiple pre-trained models can be leveraged to enhance performance across a diverse range of datasets.
@@
To understand how AutoVP would perform in a data-limited regime, we gradually and uniformly reduced the amount of training data to 50%, then 25%, then 10%, and finally 1% of each training dataset’s original size. The experimental results in Figure 3 indicate that AutoVP consistently outperformed LP across all 12 datasets, and that its relative performance was especially high in the two scenarios with the lowest data volumes, i.e., 10% and 1% data usage. The dataset-specific results can be found in Figure 8 (within Appendix B.3).
@@ 4.2 Ablation Studies of AutoVP
To address the bottleneck in hyperparameter tuning caused by FullyMap, we introduced weight initialization (WI). This allows FullyMap to initialize with a more informative mapping based on the semantic meaning of the class names. As mentioned in Section 3, AutoVP’s FullyMap consists of a linear layer, which can be characterized as $L_{\text{t}}=w\cdot L_{\text{s}}+b$ where the weight $w$ is a $K_{\text{t}}$ by $K_{\text{t}}T$ real matrix, $K_{\text{t}}$ denotes the number of target classes, and $T$ is the number of templates used in the CLIP text encoder. The weight initialization is to assign the diagonal of $w$ to be 1, and the rest of it is set to 0, resulting in $w=(\bm{I}_{K_{\text{t}}}|\bm{0})$, where $\bm{I}_{K_{\text{t}}}$ is an identity matrix of size $K_{\text{t}}$, and $\bm{0}$ is a $K_{\text{t}}\times K_{\text{t}}(T-1)$ zero matrix, indicating the connection between each target class class_name and its corresponding text prompts (“This is a photo of a [class_name]”). In Figure 6 (within Appendix A.2), we represent this concept visually. As shown in Figure 3, our experimental results demonstrate that when utilizing CLIP as the pre-trained model, FullyMap with proper WI (hatched bar) can also outperform other mapping approaches.
@@
When replicating the experimental setting as shown in Table 2 and establishing a direct connection between the fully connected output mapping layer and the CLIP image encoder without incorporating the text encoder, there was a substantial decrease in average accuracy: to 69.0% (see Figure 12, Appendix B.5). Dataset-specific accuracy drops were particularly prominent in Flowers102, Food101, and OxfordIIITPet. These outcomes suggest that label semantics play a crucial role in those datasets.
@@
We also investigated the effects on the overall performance of removing the module of visual prompts from the AutoVP pipeline while retaining the pre-trained model and output-mapping modules. When the ResNet18 model was used, leaving the visual prompts in yielded better performance than omitting them in just three out of 12 cases: i.e., the SVHN, GTSRB, and ISIC datasets (Figure 10 (b), Appendix B.4). For the remaining datasets, the inclusion of visual prompts actually led to a decline in performance. This suggests that when a relatively small source model is used for VP, a significant improvement in accuracy of the sort reported in Table 6 can primarily be attributed to fully connected layer mapping, and visual prompts may be perceived as noise. On the other hand, when the CLIP model was used, most of the datasets had positive gain values (Figure 5 (b)), indicating improved performance, when visual prompts were included. This suggests that CLIP is more suitable for VP tasks than ResNet18 is.
@@
In Appendix B.8, we also conducted an analysis from a frequency perspective (Brigham, 1988) to study the generalization of visual prompt patterns. The results highlighted the effectiveness of prompting with CLIP, harnessing low-frequency features that generalize to the target domain.
Figure 4: Models Robustness with CLIP. The comparison of accuracy drop on the CIFAR10-C dataset between AutoVP and LP.
Figure 4: Models Robustness with CLIP. The comparison of accuracy drop on the CIFAR10-C dataset between AutoVP and LP.
@@ 5 Discussions
AutoVP provides joint optimization of its multiple configurations and selects different parameters according to its target tasks. In terms of output label mapping, FullyMap exhibits superior performance in vision models, but IterMap or FreqMap appear to enhance the performance of text-image models like CLIP. In this context, weight initialization with FullyMap plays an important role in CLIP, making this option one of the more frequently chosen output-mapping strategies (Table 2). We also observed that novel designs exploiting larger image scales and mapping a larger number of source classes tended to yield enhanced performance. More information on selection preferences during hyperparameter tuning can be found in Appendix B.9.
@@
We trained AutoVP on CIFAR10 and evaluated its robustness on the corrupted dataset CIFAR10-C, which consists of 18 types of filters or noise. As shown in Figure 5, AutoVP maintained greater robustness than LP regardless of whether CLIP or ResNet18 was used as the model. Its loss of accuracy, as compared to LP, was relatively small: suggesting that AutoVP exhibits a lower degree of overfitting to the training data and possesses a higher ability to resist the impact of noise than LP does. When we investigated the differences between models, we found that using ResNet18 led to better robustness than using CLIP did (Figure 11, Appendix B.4).
@@
We also evaluated accuracy gains with ResNet18 pre-trained on ImageNet-1K (Russakovsky et al., 2015) (Figure 10, Appendix B.4) and, to assess domain differences between ImageNet-1K and other downstream datasets, we calculated the FID score (Heusel et al., 2017). The results were consistent with the cases using CLIP. In conclusion, AutoVP is suitable for datasets that exhibit more OOD characteristics than the source domain dataset.

### v2

@@ 1 Introduction
Originating in the domain of natural language processing, prompting (Gao et al., 2021; Lester et al., 2021; Shi et al., 2023) has gained considerable popularity as a parameter-efficient fine-tuning approach for adapting pre-trained models to downstream tasks. Prompting’s methodology has recently been extended to the field of computer vision, where it is termed visual prompting (VP) (Bahng et al., 2022). In its simplest form, VP can be perceived as an in-domain model-reprogramming technique (Chen, 2022). More specifically, it adjusts the inputs and outputs of a pre-trained vision model to address downstream image-classification tasks, without having to make any changes to the weights or architecture of the source model’s pre-trained backbone. As such, it stands in contrast to conventional transfer-learning methods that involve complete fine-tuning, LP (i.e., involving modifications of the trainable linear layer in the penultimate layer’s output), or zero-shot learning (Radford et al., 2021). For instance, as illustrated in Figure 1, VP adds a universal trainable data frame to the target samples at the model-input stage, and then employs a mapping function – which can be either explicitly defined or implicitly learned – to associate the source and target labels at the output stage.
@@
ILM-VP
ResNet18
CLIP
48
30
IterMap 1
45.19
78.45
@@ 2 Background and Related Work
Traditionally, to derive task-specific machine-learning models, researchers have to train or update all model parameters. But, amid the advancement of powerful foundation models, model fine-tuning and training from scratch have both become time-consuming and parameter-inefficient approaches, usually requiring large amounts of training data and storage space. To this end, VP, also known as in-domain model reprogramming, has emerged as an effective means of obtaining machine-learning models for various domain-specific tasks (Chen, 2022). A well-developed pre-trained model from a source domain can be directly used for performing tasks in the target domain with little transformation of the target data. For example, we can use an ImageNet pre-trained model to classify medical images without modifying any of its parameters (Tsai et al., 2020). On the other hand, VP, along with temperature scaling, can also be used as a post-processing calibration method to align model confidence and accuracy (Tang et al., 2022; Hsiung et al., 2023). As compared to traditional approaches such as transfer learning, model fine-tuning, and training from scratch, VP is a low-complexity and model-agnostic strategy; and it is especially suitable for low-data domains.
@@
A VP framework can generally be divided into two trainable modules, one for input transformation and the other for output transformation. These are respectively placed at the input and output ends of a pre-trained model. In the case of input transformation, previous literature has proposed various ways to generate and place visual prompts. One of the most popular such approaches is to pad a frame around the target task image and then fill it with trainable additive-input perturbation (prompts) (Tsai et al., 2020; Chen et al., 2023b; Elsayed et al., 2019; Bahng et al., 2022; Wu et al., 2022; Oh et al., 2023). Next, since the output logits of the source pre-trained model are still in the source domain, further output transformation (e.g., LM) is required to obtain the target-domain logits. One naive way of achieving this is randomly mapping (RandMap) $m$ source labels onto the target labels. Tsai et al. (2020) found that frequency-based LM (FreqMap), which constructs its LM from the source-label prediction distribution of the target-domain data, can further improve the accuracy of downstream tasks. However, Chen et al. (2023b) argued that FreqMap lacks interpretability and that its interaction with VP is difficult to measure. To address that problem, the authors proposed iterative LM (IterMap), a transformation of FreqMap that enables it to concurrently design LM and visual prompts. Yang et al. (2023), meanwhile, proposed a semantics-based LM approach that aligns source and target classes that have similar semantic meanings. And Liao et al. (2023) utilized a prototypical verbalizer to map a mask token to downstream labels, thus providing a different perspective on LM. In this paper, we follow a similar design to Bahng et al. (2022), in which visual prompts are placed around images for input transformations, and there are four mapping methods for output transformations. Further details will be presented in Section 3.
@@
Although input prompting is commonly applied directly to the target image, researchers have also developed other prompting methods, such as convolutional visual prompt (Tsai et al., 2023), which learns prompting parameters in a small convolutional structure through self-supervision tasks without knowledge of ground truths, and visual prompt tuning (Jia et al., 2022; Sohn et al., 2023), which learns prompting parameters at intermediate layers of a source model. In this paper, we focus on a pixel-level VP approach using a task-specific prompting model for each image-classification task. As such, our approach closely resembles real-world scenarios in which a pre-trained source model remains unmodified, and external variations are not introduced internally.
@@ 3 AutoVP Framework
In our implementation of AutoVP, we choose frame-shape prompts as the default prompting template. Hence, the visual prompt sizes $p$ represent the width of the frame, and its actual number of parameters is $2cp(h+w-2p)$, where $c$, $w$, and $h$ are color channels, width and height respectively. Although the input image size is determined by the source model, when fine-tuning to a downstream dataset from a source model, there is design freedom to resize the target images and use the remaining space for visual prompts. For instance, if the source model takes images with size $224\times 224$ as input, one can scale the target image size to $128\times 128$, resulting in the final visual prompt of size $p=(224-128)/2=48$. It was shown in Bahng et al. (2022) and Wu et al. (2022) that the prompt size ($p$) plays a key role in VP performance. To automate the process of optimizing image resizing scale, we design a gradient-based optimization algorithm to implement the input scaling module, which is implemented using kornia.geometry.transform() from the Kornia library (Riba et al., 2020). The transform() function integrates a range of geometric transformations for 2D images into deep learning, including differentiable image scaling. In addition to image resizing, the prompt size $p$ will also scale up or down to fill the remaining space. Furthermore, to facilitate the optimization of image resizing and avoid bad local minima, we set the default image size to 128 along with three initial scales: 0.5, 1.0, and 1.5 to optimize, and the corresponding prompt sizes $p$ are 80, 48, and 16 respectively. Consequently, the input scaling module allows AutoVP to obtain the optimal image resizing scale and prompt size ($p$).
@@
$\tilde{x}_{\text{t}}=\mathcal{P}(x_{\text{t}})=\text{InputScaling}_{p}(x_{\text{t}})+\underbrace{\mathcal{M}_{p}\odot\sigma(\delta)}_{\textrm{\text{Prompts}}}\text{.}$ (1)
@@
After applying the prompts to the resized image through the preceding stages, the prompted image is subsequently fed into the pre-trained model, which serves as a feature extractor to generate predictions in the source domain. We include four representative pre-trained models in our AutoVP framework: ResNet18 (He et al., 2016), ResNeXt101-IG (Mahajan et al., 2018), Swin-T (Liu et al., 2021), and a vision-language multi-modal model, CLIP (Radford et al., 2021) with the ViT-B/32 vision encoder backbone. Note that in AutoVP, the weights of the pre-trained classifiers are frozen and kept unchanged. The details of the models are provided in Appendix A.1.
@@
The pre-trained models predict target data to source labels, while the last mile for VP is to map predictions on the source labels to target classes. As illustrated in Figure 1, AutoVP provides four types of output mapping, and they can be generally categorized into two groups. (i) nonparametric label mapping: frequency mapping (FreqMap) and semantic mapping (SemanticMap), which are defined during the initialization of VP training and remain unchanged throughout the training process; and (ii) trainable label mapping: iterative label mapping (IterMap) and fully connected layer mapping (FullyMap). These two methods dynamically adjust the mapping based on the prompted images. In the following, we provide the overview of our four output mapping approaches, please refer to Appendix A.2 for more details.
@@
    The objective of FreqMap is to map the target label $y_{\text{t}}$ to the source label $y_{\text{s}}^{*}$, which is the most frequent label that $f_{\theta_{\text{s}}}$ classified $\tilde{\mathcal{X}_{\text{t}}}$ as. If a source class is selected as the most frequently predicted class for multiple target classes, it will be assigned to the target class that has the highest count of predictions. The general many-to-one frequency mapping algorithm is provided in Algorithm 1 in the Appendix A.2. Moreover, random label mapping (RandMap) can be viewed as a special case of FreqMap by randomly assigning a subset of source labels to a target label.
@@ 4.1 Experimental Results
We observed that AutoVP employed FullyMap as the output transformation on most datasets. We speculate that this might have been because the linear layer has more parameters and thus allows the achievement of better results. Also, when AutoVP selected initial image scales, it had a tendency to scale up those images with relatively small prompt sizes. This allowed the VP model to allocate more attention to the image itself, leading to improved overall performance. As shown in Figure 1, when ResNet18 was used as the source model, AutoVP outperformed ILM-VP by 24.8% on average. More experimental results under this setting are provided in Appendix C.1.
@@
We also allowed AutoVP to search the optimal source model for downstream tasks. The optimal settings selected by AutoVP, and a comparison of experimental results can be found in Appendix C.1. Our experimental results show that Swin-T was the pre-trained model most frequently chosen by AutoVP as most suitable, i.e., in the cases of eight of the 12 datasets. On average, this choice resulted in 0.43% better accuracy than when CLIP was utilized as the fixed pre-trained backbone. On the DTD and Flowers102 datasets, however, Swin-T’s performance was better than CLIP’s by much more: i.e., 6.80% and 3.08%, respectively. These findings highlight how multiple pre-trained models can be leveraged to enhance performance across a diverse range of datasets.
@@
To understand how AutoVP would perform in a data-limited regime, we gradually and uniformly reduced the amount of training data to 50%, then 25%, then 10%, and finally 1% of each training dataset’s original size. The experimental results in Figure 3 indicate that AutoVP consistently outperformed LP across all 12 datasets, and that its relative performance was especially high in the two scenarios with the lowest data volumes, i.e., 10% and 1% data usage. The dataset-specific results can be found in Figure C.1 (within Appendix C.2).
@@ 4.2 Ablation Studies of AutoVP
To address the bottleneck in hyperparameter tuning caused by FullyMap, we introduced weight initialization (WI). This allows FullyMap to initialize with a more informative mapping based on the semantic meaning of the class names. As mentioned in Section 3, AutoVP’s FullyMap consists of a linear layer, which can be characterized as $L_{\text{t}}=w\cdot L_{\text{s}}+b$ where the weight $w$ is a $K_{\text{t}}$ by $K_{\text{t}}T$ real matrix, $K_{\text{t}}$ denotes the number of target classes, and $T$ is the number of templates used in the CLIP text encoder. The weight initialization is to assign the diagonal of $w$ to be 1, and the rest of it is set to 0, resulting in $w=(\bm{I}_{K_{\text{t}}}|\bm{0})$, where $\bm{I}_{K_{\text{t}}}$ is an identity matrix of size $K_{\text{t}}$, and $\bm{0}$ is a $K_{\text{t}}\times K_{\text{t}}(T-1)$ zero matrix, indicating the connection between each target class class_name and its corresponding text prompts (“This is a photo of a [class_name]”). In Figure A.1 (within Appendix A.2), we represent this concept visually. As shown in Figure 3, our experimental results demonstrate that when utilizing CLIP as the pre-trained model, FullyMap with proper WI (hatched bar) can also outperform other mapping approaches.
@@
When replicating the experimental setting as shown in Table 2 and establishing a direct connection between the fully connected output mapping layer and the CLIP image encoder without incorporating the text encoder, there was a substantial decrease in average accuracy: to 69.0% (see Figure E.1, Appendix E.1). Dataset-specific accuracy drops were particularly prominent in Flowers102, Food101, and OxfordIIITPet. These outcomes suggest that label semantics play a crucial role in those datasets.
@@
We also investigated the effects on the overall performance of removing the module of visual prompts from the AutoVP pipeline while retaining the pre-trained model and output-mapping modules. When the ResNet18 model was used, leaving the visual prompts in yielded better performance than omitting them in just three out of 12 cases: i.e., the SVHN, GTSRB, and ISIC datasets (Figure C.3 (b), Appendix C.3). For the remaining datasets, the inclusion of visual prompts actually led to a decline in performance. This suggests that when a relatively small source model is used for VP, a significant improvement in accuracy of the sort reported in Table C.1 can primarily be attributed to fully connected layer mapping, and visual prompts may be perceived as noise. On the other hand, when the CLIP model was used, most of the datasets had positive gain values (Figure 5 (b)), indicating improved performance, when visual prompts were included. This suggests that CLIP is more suitable for VP tasks than ResNet18 is.
@@
In Appendix D.1, we also conducted an analysis from a frequency perspective (Brigham, 1988) to study the generalization of visual prompt patterns. The results highlighted the effectiveness of prompting with CLIP, harnessing low-frequency features that generalize to the target domain.
Figure 4: Models Robustness with CLIP. The comparison of accuracy drop on the CIFAR10-C dataset across AutoVP, ILM-VP, CLIP-VP and LP.
Figure 4: Models Robustness with CLIP. The comparison of accuracy drop on the CIFAR10-C dataset across AutoVP, ILM-VP, CLIP-VP and LP.
@@ 5 Discussions
AutoVP provides joint optimization of its multiple configurations and selects different parameters according to its target tasks. In terms of output label mapping, FullyMap exhibits superior performance in vision models, but IterMap or FreqMap appear to enhance the performance of text-image models like CLIP. In this context, weight initialization with FullyMap plays an important role in CLIP, making this option one of the more frequently chosen output-mapping strategies (Table 2). We also observed that novel designs exploiting larger image scales and mapping a larger number of source classes tended to yield enhanced performance. More information on selection preferences during hyperparameter tuning can be found in Appendix D.3.
@@
We trained AutoVP on CIFAR10 and evaluated its robustness on the corrupted dataset CIFAR10-C, which consists of 18 types of filters or noise. As shown in Figure 5, AutoVP maintained greater robustness than ILM-VP, CLIP-VP, and LP. Its loss of accuracy was relatively small: suggesting that AutoVP exhibits a lower degree of overfitting to the training data and possesses a higher ability to resist the impact of noise than the other baselines do.
@@
We also evaluated accuracy gains with ResNet18 pre-trained on ImageNet-1K (Russakovsky et al., 2015) (Figure C.3, Appendix C.3) and, to assess domain differences between ImageNet-1K and other downstream datasets, we calculated the FID score (Heusel et al., 2017). The results were consistent with the cases using CLIP. In conclusion, AutoVP is suitable for datasets that exhibit more OOD characteristics than the source domain dataset.